A mass of 4kg is first weighed on a balance at the top of a tower 25m high. The mass is then suspended from a fine wire 25m long and reweighed . Find the differences in weight. Assume the Earth is 6400km , the mass of the earth is 6*10²5kg and G is 6.67*10—¹¹Nm²kg—².
1
Expert's answer
2020-03-18T10:27:21-0400
"F=\frac{GmM}{R^2}"

"\Delta F=GmM\left(\frac{1}{R^2}-\frac{1}{(R+h)^2}\right)"

"\Delta F=( 6.67\cdot10^{-11})(4)(6\cdot10^{24})\cdot\\\cdot\left(\frac{1}{6400000^2}-\frac{1}{6400025^2}\right)=3.05\cdot10^{-4}N"